OpenStudy (anonymous):

Which of the following is the vertex of the quadratic equation y = 6x^2 - 24x + 32? (-2, -8) (-2, 8) (2, -8) (2, 8)

OpenStudy (anonymous):

so i know the vertex form is y=a(x-h)^2 +k

OpenStudy (anonymous):

idk how to turn y=6x^2 - 24 + 32 into the vertex form

OpenStudy (driftracer305):

factorize it....

OpenStudy (campbell_st):

so start this way y = 6(x^2 - 4x) + 32 what needs to be added to get the perfect square...?

OpenStudy (anonymous):

i dunno

OpenStudy (campbell_st):

well the perfect square would be y = 6(x^2 - 4x + 4) + 32 so how do you compensate..?

OpenStudy (campbell_st):

or what would I have added to create the perfect square in the brackets..?

OpenStudy (anonymous):

+4 ?

OpenStudy (campbell_st):

yep... but because the common factor of 6 is outside I really needed to add 24 6 x 4 = 24... so compensate I now subract 24 from 32 so you have \[y = 6(x^2 -4x + 4) + 32 - 24\] wich can be simplified to \[y = 6(x -2)^2 + 8\] so any ideas on the vertex...?

OpenStudy (anonymous):

(2,8)

OpenStudy (campbell_st):

yep... looks good to me.... to check you can subsitute x = 2 into the equation and you should get y = 2

OpenStudy (campbell_st):

oops y = 8